OpenStudy (anonymous):

Find the unit vector that has the same direction as the vector v = –12j

jimthompson5910 (jim_thompson5910):

given a vector v, the unit vector u that points in the same direction v is \[\Large \vec{u} = \frac{1}{|\vec{v}|}*\vec{v}\]

jimthompson5910 (jim_thompson5910):

the portion |v| is the length of vector v

OpenStudy (anonymous):

i did some research and figured it was a -v

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

-12v/12v=-v

jimthompson5910 (jim_thompson5910):

I think you meant to say -j

OpenStudy (anonymous):

or negative j sorry so many letters, -12j/12j=-j

jimthompson5910 (jim_thompson5910):

-12j = 0i + (-12j) = <0, -12>

jimthompson5910 (jim_thompson5910):

this vector is of length 12

jimthompson5910 (jim_thompson5910):

so you multiply the components of <0, -12> by 1/12

jimthompson5910 (jim_thompson5910):

to get <0,-1> which is really -j

OpenStudy (anonymous):

same answer i found just written differently

jimthompson5910 (jim_thompson5910):

yeah